How do i solve the system for x and y? I need to solve for 1/x and 1/y first though.
5/x-4/y=-180
4/x-5/y=-180

Answer by stanbon(75887)

5/x-4/y=-180
4/x-5/y=-180
----
Let 1/x = a
Let 1/y = b
---------------
5a - 4b = -180
4a - 5b = -180
-----------------------
Multiply thru 1st by 4
Multiply thru 2nd by 5
------------------------------
20a - 16b = 4*-180
20a - 25b = 5*-180
----
Subtrace and solve for "b":
9b = 180
b = 20
---
So 1/y = 20 and y = 1/20
--------------------------------
Solve for "a", then "x":
4a - 5b = -180
4a - 5*20 = -180
4a -100 = -180
4a = -80
a = -20
---
So 1/x = -20
x = -1/20
===================
Cheers,
stan H.
====================